rectangular open channel 5m wide carrying a flow of 8 m3/s. Slope of the channel bed is 0.00652. Hydraulic jump occurs in the downstream part of a channel and produced a sequent depth of 2.5m. use n=0.015.

a. Determine the depth before the jump in 


1
Expert's answer
2021-05-09T13:06:27-0400

"F^2 = \frac{Q^2T}{gA^3} \\

= \frac{8^2 \times 5}{9.81 \times (5 \times 2.5)^3} \\

= 0.0167 \\

\frac{y_2}{y_1} = \frac{1}{2}(-1 + \sqrt{1+QF^2}) \\

y_2 = 2.5 \times \frac{1}{2}(-1 + \sqrt{1 + 8 \times 0.0167}) \\

= 0.08 \;m"